OpenStudy (anonymous):

what is another way to write the expression 1/5 divided by 3

OpenStudy (sbuck98):



OpenStudy (kayders1997):

You can multiply by the reciprocal when diving you keep the first the same and flip the second number than you can multiply

OpenStudy (anonymous):

the choices are

OpenStudy (sbuck98):

if you solve it that's what you get^^

OpenStudy (anonymous):

1/5 divided by 1/3

OpenStudy (anonymous):

1/5+3

OpenStudy (anonymous):

1/5+1/3

OpenStudy (anonymous):

1/5 times 1/3

OpenStudy (kayders1997):

Pretend you have 3/1 flip that fraction because it is the second fraction diving fractions don't ask when flip the secon number and multiply

OpenStudy (jdoe0001):

any number, any whatsoever, can be written as "over 1" thus 3 is really 3/1 thus \(\bf \cfrac{\frac{1}{5}}{3}\iff \cfrac{\frac{1}{5}}{\frac{3}{1}}\implies \cfrac{1}{5}\cdot \cfrac{1}{3}\implies \cfrac{1}{15}\)

OpenStudy (kayders1997):

Keep first number the same the same so you will have 1/5 flip second fraction 3/1 to 1/3 than multiply so 1/5 times 1/3

OpenStudy (jdoe0001):

and recall that \(\large \cfrac{\frac{a}{b}}{\frac{c}{{\color{blue}{ d}}}}\implies \cfrac{a}{b}\cdot \cfrac{{\color{blue}{ d}}}{c}\)

OpenStudy (anonymous):

thank @kayders1997 and @jdoe0001

OpenStudy (kayders1997):

No problem Jaredalm :)